# Automatic Database Configuration Debugging using Retrieval-Augmented Language Models

### v1

@@ Abstract.
conference: 2020 International Conference on Management of Data; June 14–19, 2020; Portland, OR
@@ 1. Introduction
Clearly, the above process is challenging and time-consuming, as it requires the DBA to be highly experienced in configuration debugging and have a comprehensive understanding of the DBMS’s internal working mechanisms. This raises a crucial question: can we replace the DBA with an LLM-based agent to enable automatic DBMS configuration debugging?
@@
(3) Extensive experiments. Experiments using real-world DBMS configuration debugging datasets show that Andromeda significantly outperforms existing alternatives. https://github.com/axzs0987/Andromeda/.
@@ 2.1. Problem Formalization
Configurable Knobs. Although there exist different ways to configure the DBMS (Dias et al., 2005a; Yoon et al., 2016a), this paper focuses on diagnosing and tuning the DBMS configurable knobs (or knobs for short). Specifically, modern DBMSs usually have hundreds of configurable knobs that control their runtime behavior, e.g., $1012$ knobs (such as wait_timeout, unique_checks) in MySQL 8.0 and more than $300$ knobs (such as trace_locks) in PostgreSQL 12.0, which play crucial roles in DBMS tuning (Fekry et al., 2020; Aken et al., 2017).
@@ 2.2. Related Work
However, these studies are not designed for taking NL questions from database users as input and recommend configurations corresponding to the issues, which is a key feature of our Andromeda framework. Moreover, they are constructed relying on pre-defined optimization objectives (such as runtime), making them difficult to address runtime errors and user-specific configurations.
@@ 3. The Andromeda Framework
As shown in Figure 3, Andromeda employs a retrieval-augmented generation (RAG) strategy that consists of offline and online stages.
@@
  • Troubleshooting Manuals ${M}$ consists of documents for DBMS manuals, e.g., MySQL Manual (mso, [n. d.]) and PostgreSQL Manual (pg_, [n. d.]), which describe how to troubleshoot configuration issues. Specifically, we first chunk each document in the manuals into sentences $\{m_{i}\}$, and then use regular-expression rules to match each sentence $m_{i}$ with the recommended configurations $K_{i}$. Consider the example in Figure 2 again. A document from MySQL Manual 10.5.5: Bulk Data Loading for InnoDB Tables contains a sentence “When importing data into InnoDB, turn off autocommit mode …” with its corresponding configuration $\texttt{autocommit}=0$. In such a way, we form a set of troubleshooting manuals ${M}=\{\langle m_{i},K_{i}\rangle\}$.
Given document sets ${Q}$ and ${M}$, Andromeda utilizes a Document Encoder to map each document (either historical question $q^{\tt H}$ or manual instance $m$) into a vector-based embedding representation. Then, it stores learned document embeddings in a vector database, which will be further used for online document retrieval.
The key challenge here is that the semantics of documents from different sources may be heterogeneous. Consider the examples of historical questions and troubleshooting manuals in Figure 2 again: historical questions typically describe “what is wrong”, while manuals focus on “How to fix it”. Thus, it is challenging to develop a unified document encoder that understands the semantics of different types of documents for effective document retrieval. To address this challenge, we propose an effective document encoding mechanism that addresses the semantic heterogeneity among different sources. We propose a contrastive learning-based approach that aligns documents from different sources into a unified embedding space, and enables accurate document retrieval on the space. We also design a data synthesis strategy to tackle the difficulty of limited training data. We will provide more details in Section 4.
@@
Step 2: Telemetry Analysis. Inspired by the way a DBA performs configuration debugging, Andromeda also needs to know what is happening in the DBMS $D$. To this end, we propose to extract key features from the telemetry data, i.e., time-series data of performance metrics (e.g., CPU utilization). Formally, we define the telemetry data as a set ${S}=\{s_{i},\{t^{i}_{j}\}\}$ of time-series performance metrics, where each instance consists of a performance metric $s_{i}$ and a collection of timestamps $\{t^{i}_{j}\}$. Note that not all performance metrics or all timestamps of a particular metric may be relevant to the user’s question $q$. Thus, Andromeda utilizes a Telemetry Analyzer to detect “troublesome” telemetries $\{s_{i}\}$ with the particular timestamps when a performance issue occurs (e.g., the duration of full CPU usages). Moreover, the Telemetry Analyzer also suggests the related knobs that need adjustment. Consider the example in Figure 2. Andromeda considers telemetries of DBMS $D$ and detects three troublesome performance metrics, e.g., innodb_log_write_requests, and suggests the related knobs that need adjustment, e.g., innodb_log_file_size.
@@
Step 3: Configuration Reasoning. With the above useful and contextual information, Andromeda employs a Configuration Reasoner that leverages an LLM (e.g., GPT-4 or Llama 3) to diagnose improper DBMS configurations and generate accurate troubleshooting recommendations. Specifically, it generates a prompt that consists of three components: (1) task instruction, (2) retrieved data, including historical questions, troubleshooting manuals and relevant telemetries, and (3) user’s NL question $q$. Figure 2 provides an example of the prompting mechanism (the central part of the figure). After that, Andromeda prompts an LLM to generate the configuration recommendations (as shown in the right part of Figure 2). In particular, we utilize a two-phase prompting strategy that respectively (1) identify a subset $K\subseteq\mathbb{K}$ of knobs relevant to the corresponding configuration issues in $q$, and (2) recommend a proper value $v\in\mathbb{D}_{k}$ for tuning each relevant knob $k\in K$, which makes sense because (i) Using RAG, LLM can obtain knowledge about the knobs. For example, LLM can set autocommit=0 with manual, and it can set innodb_log_file_size = 256 if the similar question sets the same value. (ii) LLMs may learn some database related knowledge in pre-training or fine-tuning process, which is helpful for knob value decisions. Note that there could be multiple ways, such as different tuning knobs, to fix the same issue. Andromeda uses the LLM to reason about the various tuning knobs and select the most appropriate set. Specifically, Andromeda generates a prompt that includes all the retrieved data and then calls the LLM to recommend the configurations.
@@
To use Andromeda flexibly, users can selectively utilize the sources (for instance, if telemetry is unavailable, they can only consider documents). Moreover, users can either use only Step 1 to identify the correct subset of knobs or perform an end-to-end process to predict the knob values.
@@
The key challenge in the online stage is how to detect troublesome telemetries. Specifically, there are many types of telemetries in a DBMS, and only a small number of telemetries are relevant to the question. Moreover, even given a relevant telemetry, only a small proportion of timestamps are useful. Thus, it is non-trivial to detect such troublesome telemetries that are relevant to a user’s NL question. To address the challenge, we introduce a telemetry analysis method to identify troublesome telemetries that are relevant to a user’s NL question. The method first utilizes a seasonal-trend decomposition mechanism to find troublesome telemetries, and then judiciously selects the telemetries relevant to a user’s question. Please refer to Section 5 for more details.
@@
Discussion on Generalizability. Andromeda can be easily generalized to other DBMSes, as the algorithms for training data generation, model training, and evaluation are adaptable to new DBMSes. When applying Andromeda to a new DBMS, all that is required is the specification of the knobs and the manuals of the DBMS. Andromeda can then use these inputs to automatically generate augmented data and train a new model during the offline stage, while supporting evaluation on the DBMS in the online stage.
@@ 4. Multi-Source Document Retrieval
Secondly, constructing such an encoder $f_{\tt enc}$ that aligns documents from different sources requires a substantial amount of training data. However, in practice, there is often an insufficiency in both quantity and quality of the training data. To address the challenge, we propose a logic-based task decomposition method that uses LLMs to perform logical reasoning with the knowledge from manuals, thereby augmenting high-quality training data.
@@ 4.1. Document Representation Learning
where $d^{+}\in D^{+}$ is positive document, while $D^{-}$ are negative documents. The basic idea of the above InfoNCE loss brings positive retrieval texts closer in the embedding space while pushing negative retrieval texts further apart, as shown in Figure 4 (b).
@@ 4.2. Training Data Augmentation
For better training $f_{\tt enc}$, we need a sufficient amount of training data, i.e., a large number of target queries ($q^{*}$) with ground-truth knob configurations. However, it is time- and effort-consuming to collect a large amount of high-quality training data, which hinders the effectiveness of document representation learning. To solve this problem, we introduce a novel data synthesis mechanism that utilizes an LLM to automatically generate NL questions with corresponding knob configurations from DBMS manuals.
(a) An example of logic chain.
(b) Data synthesis using a logic chain strategy.
@@
Our Logic Chain Strategy. To address the above limitation, we introduce a logic-chain based strategy for effective data synthesis. The basic idea is to generate realistic questions by simulating a real user’s experience when encountering a DBMS configuration issue. Specifically, we consider the natural logical reasoning chain from issues to the solutions (e.g., recommended knob configurations):
@@
Intuitively, if we traverse the logic chain reversely, we can inject domain knowledge about the logical relationships between knobs and issues into the LLM. Moreover, as the functions can be quite diverse, given a knob, we can generate a wide range of possible questions, which is helpful to improve the diversity of training data.
Example 0 (Our Logic Chain based Synthesis Strategy.).
Consider the manual of max_allowed_packet in the previous example again. We can extract diverse knob functions from the manual, and then we can infer that function “set an upper limit on the output length” can resolve issue “fail to output due to excessive content length”, which is triggered by the behavior “executes a repeat command”. We provide such a logic chain to the LLM, and obtain a more realistic question “I want to connect to the remote SQL server to simulate traffic using the REPEAT command in MySQL. And I executed SQL: REPEAT(’A’, 26214400). The connection is successful, but I encounter an error that returns NULL. How can I solve this for MySQL?”
Based on the above idea, we utilize an LLM with a logic-chain prompting to generate questions. Specifically, we decompose the data synthesis task into several sub-tasks. Each sub-task aims to generate details for each edge in the logic chain using the LLM, as shown in Figure 5(b). For instance, for “knob->function” in the logic chain, we can construct the sub-task with question “What are the functions of max_allowed_packet?”, which can be answered as spec_func by the LLM. Next, for “function->Anomaly”, the question can be “What anomaly can be solved with the spec_func of max_allowed_packet”. When the LLM provides answers for each edge in the complete logic chain, we can integrate all the answers to identify a specific configuration debugging question.
@@ 5. Telemetry Data Analysis
We introduce a telemetry data analysis approach consisting of anomalous telemetry detection, telemetry-to-text and relevant telemetry selection. We develop techniques to address the following two challenges in the above three steps. The first challenge is detecting anomalous telemetries from a huge amount of time-series data, as it is not easy to identify what is anomalous. To address this, we first use a seasonal-trend decomposition method to decompose time series data and detect anomalous telemetries using statistics on the residual. Secondly, there may exist hundreds of telemetries, many of which may not related to our NL question $q$. To tackle this challenge, we propose a telemetry-to-text mechanism that enables to select telemetries relevant to $q$.
@@ 5.1. Anomalous Telemetry Detection
where $Seasonal_{s_{i}}$ represents the captured repeating seasonal patterns by smoothing the data within each period. $Trend_{s_{i}}$ is the long-term trend captured in the time series using local regression smoothing (LOESS), and $\epsilon$ is the unexplained random part of the time series after removing the trend and seasonal components. With STL, we can decompose seasonal and trends (normal variations), and assess whether the value on each timestamp is anomalous or not by analyzing the residual data $\epsilon_{si}$. After removing the effects of seasonal and trend components, it is easier to detect anomalous deviations for $s_{i}$ using statistic-based methods.
@@
The median absolute deviation (MAD) is the median of the absolute deviations from the median of all data points in $\lambda_{s_{i}}$, denoted as:
@@ 5.2. Telemetry-to-Text and Telemetry Selection
Due to the input-length limitations of LLM and the possibility that irrelevant metrics may mislead the LLM, we need to select useful anomalous telemetries from a large set as prompts to assist LLM to perform reasoning of DBMS diagnosis. Intuitively, we need to select metrics that are relevant to the user’s question $q$. To achieve this goal, we rank the anomalous telemetry descriptions on the basis of the semantic similarity between the descriptions and the user’s NL question (e.g., using Sentence-Bert). Finally, we select the top-$K$ anomalous descriptions as the final prompts.
@@
As shown in Figure 6, given a user’s NL question about the slow execution time of “INSERT” statement, we monitor the temporal performance $T_{S}$ as a temporal telemetry data with hundreds of metrics about CPU, Memory and MySQL system. For time-series data $T_{s_{i}}$ of each metric $s_{i}$, such as Memory_free_total recording the remaining memory, we iteratively detect the most anomalous value in $T_{s_{i}}$. In each iteration, we use STL to decompose $T_{s_{i}}$ as seasonal, trend and residual. For residual, we used the mean absolute deviation (MAD) to assess the severity of the anomaly at each timestamp $j$ and collect them. After all anomalous deviation collected (Memory_free_total, MySQL_innodb_log_wirte_requests and others), we convert them into text and select relevant telemetries using semantic similarity with question $q$. Finally, we concatenate the generated text as the result of telemetry data analysis (the right part of the figure).

### v2

@@ Abstract.
copyright: acmlicensedjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: N1 (SIGMOD)article: 013publicationmonth: 1price: 15.00doi: 10.1145/3709663
@@ 1. Introduction
Clearly, the above process is challenging and time-consuming, as it requires the DBA to be highly experienced in configuration debugging and to have a comprehensive understanding of the DBMS’s internal working mechanisms. This raises a crucial question: Can we replace the DBA with an LLM-based agent to enable automatic DBMS configuration debugging?
@@
(3) Extensive experiments. Experiments using real-world DBMS configuration debugging datasets show that Andromeda significantly outperforms existing alternatives. We make the code of Andromeda and the datasets in our experiments public on a Github repo [^1] https://github.com/ruc-datalab/Andromeda/.
@@ 2.1. Problem Formalization
Configurable Knobs. Although there exist different ways to configure the DBMS (Dias et al., 2005a; Yoon et al., 2016a), this paper focuses on diagnosing and tuning the DBMS configurable knobs (or knobs for short). Specifically, modern DBMSs usually have hundreds of configurable knobs that control their runtime behavior, e.g., $1,012$ knobs (such as wait_timeout, unique_checks) in MySQL 8.0 and more than $300$ knobs (such as trace_locks) in PostgreSQL 12.0, which play crucial roles in DBMS tuning (Fekry et al., 2020; Aken et al., 2017).
@@ 2.2. Related Work
However, these studies are not designed for taking NL questions from database users as input and recommend configurations corresponding to the issues, which is a key feature of our Andromeda framework. Moreover, they are constructed by relying on pre-defined optimization objectives (such as runtime), making them difficult to address runtime errors and user-specific configurations.
@@ 3. The Andromeda Framework
As shown in Figure 3, Andromeda employs a retrieval-augmented generation (RAG) strategy consisting of offline and online stages.
@@
  • Troubleshooting Manuals ${M}$ consists of documents for DBMS manuals, e.g., MySQL Manual (mso, [n. d.]) and PostgreSQL Manual (pg_, [n. d.]), which describe how to troubleshoot configuration issues. Specifically, we first divide each document in the manuals into sentences $\{m_{i}\}$, and then use regular expression rules to match each sentence $m_{i}$ with the recommended configurations $K_{i}$. Consider the example in Figure 2 again. A document from MySQL Manual 10.5.5: Bulk Data Loading for InnoDB Tables contains a sentence “When importing data into InnoDB, turn off autocommit mode …” with its corresponding configuration $\texttt{autocommit}=0$. In this way, we form a set of troubleshooting manuals ${M}=\{\langle m_{i},K_{i}\rangle\}$.
Given document sets ${Q}$ and ${M}$, Andromeda uses a Document Encoder to map each document (either the historical question $q^{\tt H}$ or manual instance $m$) into a vector-based embedding representation. Then, it stores the learned document embeddings in a vector database, which will be further used for online document retrieval.
The key challenge here is that the semantics of documents from different sources may be heterogeneous. Consider the examples of historical questions and troubleshooting manuals in Figure 2 again: historical questions typically describe “what is wrong”, while manuals focus on “How to fix it”. Thus, it is challenging to develop a unified document encoder that understands the semantics of different types of documents for effective document retrieval. To address this challenge, we propose an effective document encoding mechanism that addresses semantic heterogeneity among different sources. We propose a contrastive learning-based approach that aligns documents from different sources into a unified embedding space, and enables accurate document retrieval on the space. We also design a data synthesis strategy to address the difficulty of limited training data. We will provide more details in Section 4.
@@
Step 2: Telemetry Analysis. Inspired by the way a DBA performs configuration debugging, Andromeda also needs to know what is happening in the DBMS $D$. To this end, we propose to extract key features from the telemetry data, i.e., time-series data of performance metrics (e.g., CPU utilization). Formally, we define the telemetry data as a set ${S}=\{s_{i},\{t^{i}_{j}\}\}$ of time-series performance metrics, where each instance consists of a performance metric $s_{i}$ and a collection of timestamps $\{t^{i}_{j}\}$. Note that not all performance metrics or all timestamps of a particular metric may be relevant to the user’s question $q$. Thus, Andromeda utilizes a Telemetry Analyzer to detect “troublesome” telemetries $\{s_{i}\}$ with the particular timestamps when a performance issue occurs (e.g., the duration of full CPU usages). Moreover, Telemetry Analyzer also suggests the related knobs that need adjustment. Consider the example in Figure 2. Andromeda considers telemetries of DBMS $D$ and detects three troublesome performance metrics, e.g., innodb_log_write_requests, and suggests the related knobs that need adjustment, e.g., innodb_log_file_size.
@@
Step 3: Configuration Reasoning. With the above useful and contextual information, Andromeda employs a Configuration Reasoner that leverages an LLM (e.g., GPT-4 or Llama 3) to diagnose improper DBMS configurations and generate accurate troubleshooting recommendations. Specifically, it generates a prompt that consists of three components: (1) task instruction, (2) retrieved data, including historical questions, troubleshooting manuals and relevant telemetries, and (3) user’s NL question $q$. Figure 2 provides an example of the prompting mechanism (the central part of the figure). After that, Andromeda prompts an LLM to generate configuration recommendations (as shown in the right part of Figure 2). In particular, we utilize a two-phase prompting strategy that respectively (1) identify a subset $K\subseteq\mathbb{K}$ of knobs relevant to the corresponding configuration issues in $q$, and (2) recommend a proper value $v\in\mathbb{D}_{k}$ for tuning each relevant knob $k\in K$, which makes sense because (i) Using RAG, LLM can obtain knowledge about the knobs. For example, LLM can set autocommit=0 with manual, and it can set innodb_log_file_size = 256 if the similar question sets the same value. (ii) LLMs may learn some database related knowledge in pre-training or fine-tuning process, which is helpful for knob value decisions. Note that there could be multiple ways, such as different tuning knobs, to fix the same issue. Andromeda uses the LLM to reason about the various tuning knobs and select the most appropriate set. Specifically, Andromeda generates a prompt that includes all the retrieved data and then calls the LLM to recommend the configurations.
@@
To use Andromeda flexibly, users can selectively utilize the sources (for instance, if telemetry is unavailable, they can only consider documents). Moreover, users can use only Step 1 to identify the correct subset of knobs or perform an end-to-end process to predict the knob values.
@@
The key challenge on the online stage is how to detect troublesome telemetries. Specifically, there are many types of telemetries in a DBMS, and only a small number of telemetries are relevant to the question. Moreover, even given a relevant telemetry, only a small proportion of timestamps are useful. Thus, it is non-trivial to detect such troublesome telemetries that are relevant to a user’s NL question. To address the challenge, we introduce a telemetry analysis method to identify troublesome telemetries that are relevant to a user’s NL question. The method first utilizes a seasonal-trend decomposition mechanism to find troublesome telemetries, and then judiciously selects the telemetries relevant to a user’s question. Please refer to Section 5 for more details.
@@
Discussion on Generalizability. Andromeda can be easily generalized to other DBMSes, as the algorithms for training data generation, model training, and evaluation are adaptable to new DBMSes. When applying Andromeda to a new DBMS, all that is required is the specification of the knobs and the DBMS manual. Andromeda can then use these inputs to automatically generate the augmented data and train a new model during the offline stage, while supporting the evaluation of the DBMS in the online stage.
@@ 4. Multi-Source Document Retrieval
Secondly, constructing such an encoder $f_{\tt enc}$ that aligns documents from different sources requires a substantial amount of training data. However, in practice, there is often an insufficiency in both quantity and quality of the training data. To address the challenge, we propose a logic-chain based task decomposition method that uses LLMs to perform logical reasoning with the knowledge from manuals, thereby augmenting high-quality training data.
@@ 4.1. Document Representation Learning
where $d^{+}$ is positive document, while $d^{-}\in D^{-}$ are negative documents. The basic idea of the above InfoNCE loss brings positive retrieval texts closer in the embedding space while pushing negative retrieval texts farther apart, as shown in Figure 4 (b).
@@ 4.2. Training Data Augmentation
For better training of $f_{\tt enc}$, we need a sufficient amount of training data, i.e., a large number of target queries ($q^{*}$) with ground-truth knob configurations. However, it is time- and effort-consuming to collect a large amount of high-quality training data, which hinders the effectiveness of document representation learning. To solve this problem, we introduce a novel data synthesis mechanism that utilizes an LLM to automatically generate NL questions with corresponding knob configurations from DBMS manuals.
(a) An example of logic-chain.
(b) Data synthesis using a logic-chain strategy.
@@
Our Logic-Chain Strategy. To address the above limitation, we introduce a logic-chain based strategy for effective data synthesis. The basic idea is to generate realistic questions by simulating a real user’s experience when encountering a DBMS configuration issue. Specifically, we consider the natural logical reasoning chain from issues to the solutions (e.g., recommended knob configurations):
@@
Intuitively, if we traverse the logic-chain reversely, we can inject domain knowledge about the logical relationships between knobs and issues into the LLM. Moreover, as the functions can be quite diverse, given a knob, we can generate a wide range of possible questions, which is helpful to improve the diversity of training data.
Example 0 (Our Logic-Chain based Synthesis Strategy.).
Consider the manual of max_allowed_packet in the previous example again. We can extract diverse knob functions from the manual, and then we can infer that function “set an upper limit on the output length” can resolve issue “fail to output due to excessive content length”, which is triggered by the behavior “executes a repeat command”. We provide such a logic-chain to the LLM, and obtain a more realistic question “I want to connect to the remote SQL server to simulate traffic using the REPEAT command in MySQL. And I executed SQL: REPEAT(’A’, 26214400). The connection is successful, but I encounter an error that returns NULL. How can I solve this for MySQL?”
Based on the above idea, we utilize an LLM with a logic-chain prompting to generate questions. Specifically, we decompose the data synthesis task into several sub-tasks. Each sub-task aims to generate details for each edge in the logic-chain using the LLM, as shown in Figure 5(b). For instance, for “knob->function” in the logic-chain, we can construct the sub-task with question “What are the functions of max_allowed_packet?”, which can be answered as spec_func by the LLM. Next, for “function->Anomaly”, the question can be “What anomaly can be solved with the spec_func of max_allowed_packet”. When the LLM provides answers for each edge in the complete logic-chain, we can integrate all the answers to identify a specific configuration debugging question.
@@ 5. Telemetry Data Analysis
We introduce a telemetry data analysis approach consisting of anomalous telemetry detection, telemetry-to-text and relevant telemetry selection. We develop techniques to address the following two challenges in the above three steps. The first challenge is detecting anomalous telemetries from a huge amount of time-series data, as it is not easy to identify what is anomalous. To address this, we first use a seasonal-trend decomposition method to decompose time series data and detect anomalous telemetries using statistics on the residual. Secondly, there may exist hundreds of telemetries, many of which may not be related to our NL question $q$. To tackle this challenge, we propose a telemetry-to-text mechanism that enables us to select telemetries relevant to $q$.
@@ 5.1. Anomalous Telemetry Detection
where $Seasonal_{s_{i}}$ represents the captured repeating seasonal patterns by smoothing the data within each period. $Trend_{s_{i}}$ is the long-term trend captured in the time series using local regression smoothing (LOESS), and $\epsilon$ is the unexplained random part of the time series after removing the trend and seasonal components. With STL, we can decompose seasonal and trends (normal variations), and assess whether the value on each timestamp is anomalous by analyzing the residual data $\epsilon_{si}$. After removing the effects of seasonal and trend components, it is easier to detect anomalous deviations for $s_{i}$ using statistic-based methods.
@@
The median absolute deviation (MAD) is the median of the absolute deviations from the median of all data points in $\lambda_{s_{i}}$, denoted by:
@@ 5.2. Telemetry-to-Text and Telemetry Selection
Due to the input-length limitations of LLM and the possibility that irrelevant metrics may mislead the LLM, we need to select useful anomalous telemetries from a large set as prompts to assist LLM to perform reasoning of DBMS diagnosis. Intuitively, we need to select metrics that are relevant to the user’s question $q$. To achieve this goal, we rank the anomalous telemetry descriptions on the basis of the semantic similarity between the descriptions and the user’s NL question (e.g., using Sentence-BERT (Reimers and Gurevych, 2019b)). Finally, we select the top-$K$ anomalous descriptions as the final prompts.
@@
As shown in Figure 6, given a user’s NL question about the slow execution time of “INSERT” statement, we monitor the temporal performance $T_{S}$ as a temporal telemetry data with hundreds of metrics about CPU, Memory and MySQL DBMS. For time-series data $T_{s_{i}}$ of each metric $s_{i}$, such as Memory_free_total recording the remaining memory, we iteratively detect the most anomalous value in $T_{s_{i}}$. In each iteration, we use STL to decompose $T_{s_{i}}$ as seasonal, trend and residual. For residual, we use the mean absolute deviation (MAD) to assess the severity of the anomaly at each timestamp $j$ and collect them. After all anomalous deviation collected (Memory_free_total, MySQL_innodb_log_wirte_requests and others), we convert them into text and select relevant telemetries using semantic similarity with question $q$. Finally, we concatenate the generated text as the result of telemetry data analysis (the right part of the figure).